Check my workCheck My Work button is now enabled5

Item13

Item 13 2 points

Fisk Corporation is trying to improve its inventory control system and has installed an online computer at its retail stores. Fisk anticipates sales of 98,000 units per year, an ordering cost of $12 per order, and carrying costs of $1.20 per unit.

a. What is the economic ordering quantity?



b. How many orders will be placed during the year?



c. What will the average inventory be?



d. What is the total cost of ordering and carrying inventory?
Answer #1

a) Economic order quantity (EOQ) = 1400 units

Explanation:-

EOQ = square root of [(2 × units × ordering cost/carrying cost)]

EOQ= square root of (2× 98,000 × $12/$1.2)

EOQ= square root of ($2,352,000/$1.2)

EOQ= square root of (1,960,000)

EOQ= 1400

b)No. of orders = 70 orders

Explanation:-

Units/ EOQ

98,000/1400 = 70 orders

c)Average inventory = 700 units

Explanation:-

50% of EOQ

50% × 1400 =700

d) Total cost = $1,680

Explanation:-

Ordering cost = total no. Of order × order cost per unit

Ordering cost = 70 × $12 = $840

Carrying cost = EOQ × carrying cost/2

Carrying cost = 1400 ×1.2/2 = $840

Total cost of ordering and carrying cost = ordering cost + carrying cost

Total cost of ordering and carrying cost = $840 +$840 = $1,680